# VA-GS: Enhancing the Geometric Representation of Gaussian Splatting via View Alignment

### v1

@@ 5 Experiments
Evaluation Protocols. We evaluate our surface reconstruction performance on the DTU [18] and Tanks and Temples (TNT) [22] datasets. Following prior works [16, 56, 4, 57], we use 15 scenes from the DTU dataset and 6 scenes from the TNT dataset for evaluation. Depth maps are rendered for all training views, and a TSDF [7] is constructed for mesh extraction. For novel view synthesis, we use the Mip-NeRF 360 dataset [2], which contains large-scale indoor and outdoor scenes with complex lighting and fine-grained geometric details. Following 3DGS [21], one out of every eight images is used for evaluation, while the remaining seven are used for training. We employ COLMAP [36] to generate a sparse point cloud from the original dataset images for initializing the 3D Gaussians. All images are downsampled to a lower resolution to facilitate training. Following established protocols [16, 56, 4, 57], we report Chamfer distance for surface reconstruction on the DTU dataset and F1-score for the TNT dataset. For novel view synthesis, we evaluate using three widely adopted image quality metrics: PSNR, SSIM, and LPIPS.
Implementation Details. Our overall pipeline, training strategy, and hyperparameter settings generally follow 3DGS [21]. We set the number of source views to N=3N\!=\!3, the threshold in $\mathcal{L}_{ns}$ to $\tau\!=\!0.01$, and the patch size in $\mathcal{L}_{p}$ to $7\!\times\!7$. The loss weight factors are set as follows: $\beta_{1}\!=\!0.2$, $\beta_{2}\!=\!0.03$, $\lambda_{1}\!=\!0.015$, $\lambda_{2}\!=\!0.3$, $\lambda_{3}\!=\!0.15$, and $\lambda_{4}\!=\!1.0$. The model is trained for 20,000 iterations for surface reconstruction and 30,000 iterations for novel view synthesis. We first pretrain the model using only the color loss for 7,000 steps to obtain a coarse geometric initialization, which provides a stable foundation for subsequent geometry refinement. Then, we incorporate our image edge item and normal-based geometry alignment into the training. To further refine geometry, we sequentially apply our multi-view photometric alignment for 8,000 iterations, followed by 5,000 iterations of multi-view feature alignment. For novel view synthesis, we continue training for an additional 10,000 steps to optimize rendering quality. All experiments are conducted on a single NVIDIA RTX 4090 GPU.
@@ 5.1 Performance Evaluation
Comparisons on DTU. We first compare our method with state-of-the-art implicit and explicit surface reconstruction approaches on the DTU dataset [18]. Following standard protocol, reconstructions are clipped using the provided mask, and evaluations are performed only on foreground objects, as the ground truth point clouds exclude background regions. As shown in Table 1, our method achieves the lowest average Chamfer distance and ranks best across most scenes. Compared to implicit approaches such as NeuS [44] and Neuralangelo [27], our method delivers significantly better reconstruction accuracy while being much more efficient in terms of runtime. It is worth noting that most implicit methods [44, 27] only reconstruct foreground geometry, whereas our approach can produce detailed and complete meshes, including background regions, which is an essential feature for mesh-based rendering. Although our method is slightly slower than 3DGS [21] and 2DGS [16] due to the use of multi-view alignment, it achieves significant improvements in reconstruction quality over these earlier Gaussian-based methods.
@@
Comparisons on TNT. We further evaluate our method on the TNT dataset [22], comparing it against both implicit and explicit surface reconstruction baselines. Since the ground-truth point clouds do not include background regions, the evaluation is restricted to foreground objects. As shown in Table 2, our method achieves the best reconstruction performance among all competing approaches, including both implicit and explicit methods. Notably, while several Gaussian-based methods require less optimization time, they tend to produce results with much lower accuracy. In contrast, our method reaches a better balance between efficiency and reconstruction quality. For example, GS-Pull [60] only reconstructs foreground objects and often generates overly smooth surfaces. Fig. 3 provides a qualitative comparison. Our method produces more accurate and detailed reconstructions for both foreground and background regions. It also effectively mitigates the impact of shadows, whereas baseline methods often yield noisy meshes or fail to capture geometric details. The use of geometry, photometric, and feature-based alignment from multiple views provides strong guidance, enabling the Gaussian primitives to converge more accurately to the true surface geometry.
@@
Comparisons on Mip-NeRF 360. We also evaluate our approach on the Mip-NeRF 360 dataset [2] for novel view synthesis. Table 3 reports quantitative comparisons against state-of-the-art Gaussian-based and other neural rendering baselines. Our method outperforms competitors on most metrics, demonstrating superior image fitting and generalization to unseen viewpoints. This evidences that our enhanced geometry representation yields higher visual fidelity. Notably, the Mip-NeRF 360 itself achieves the highest average PSNR on indoor scenes but lags on SSIM and LPIPS. Among Gaussian-based methods, 2DGS [16], SuGaR [14], and GS-Pull [60] perform worse than vanilla 3DGS [21], suggesting that their planar Gaussian constraints degrade performance in complex environments. Our ablation results in Table 4 further confirm that flattening 3D Gaussians into planar Gaussian disks is ineffective for our framework. Our method preserves the full 3D Gaussian representation and delivers high-quality surfaces without sacrificing novel-view rendering quality. Fig. 4 provides a qualitative comparison of reconstructed meshes. Consistent with our observations on the TNT dataset, our method recovers more accurate and complete surfaces in both foreground and background regions, whereas other methods suffer from noise, oversmoothing, or missing details, especially in challenging indoor scenes.
@@ 5.2 Ablation Studies
(6) Scale regularization: The scaling matrix $\bm{S}$ represents the stretching of a spherical Gaussian along the three axes. Different from previous works [4, 6, 60], incorporating the widely used scale penalty into our method to flatten the 3D Gaussian disks provides no performance gains, and even degrades novel-view rendering quality on the Mip-NeRF 360 dataset.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China (62402401), the Sichuan Provincial Natural Science Foundation of China (2025ZNSFSC1462) and the Fundamental Research Funds for the Central Universities.
@@ Appendix A Baseline Methods
We compare our approach with state-of-the-art surface reconstruction and novel view synthesis methods based on 3D Gaussian Splatting, such as 3DGS [21], 2DGS [16], GOF [56], and PGSR [4]. In addition, we evaluate against implicit NeRF-based methods, such as NeuS [44] and Neuralangelo [27], which utilize Signed Distance Functions (SDFs) to represent scenes and convert them into opacity fields for volume rendering via ray tracing.
@@ B.1 Comparisons on the Deep Blending Dataset
In order to further test the performance of the algorithm on different real-world data, we evaluate both novel view synthesis and surface reconstruction on two commonly used indoor scenes, Dr Johnson and Playroom, from the Deep Blending dataset [15]. Table 5 presents quantitative comparisons against state-of-the-art Gaussian-based methods on novel view synthesis. All results, unless otherwise noted, are obtained from our own re-implementations using the official code released by the respective authors. For 3DGS [21], we use the pretrained models provided by the authors. For SuGaR [14], we report numbers directly from the original paper to avoid potential inconsistencies. Our method consistently outperforms all baselines across all metrics, demonstrating superior rendering quality and generalization to unseen viewpoints. These improvements can be attributed to our enhanced geometric representation, which yields better visual fidelity. Consistent with observations in the evaluation on the Mip-NeRF 360 dataset [2] presented in the main paper, methods such as 2DGS [16], SuGaR [14], and GS-Pull [60] underperform compared to vanilla 3DGS [21], suggesting that imposing planar Gaussian constraints may hinder performance in complex scenes.
@@ B.2 More Experimental Results
Novel view synthesis. To provide a more comprehensive evaluation of novel view synthesis, we report per-scene quality metrics (PSNR, SSIM, and LPIPS) for Gaussian-based methods on the Mip-NeRF 360 dataset [2], as shown in Table 6. All results are obtained from our own runs using the official implementations of prior methods, except for 3DGS, where we use the pretrained models provided by the authors to avoid training inconsistencies. Our method consistently achieves the highest PSNR and SSIM scores on most scenes and has the best average scores. Notably, it yields significant improvements in LPIPS across all scenes, highlighting its ability to capture high-frequency details and produce perceptually superior renderings.
@@
Surface reconstruction. Fig. 7 presents additional qualitative comparisons of reconstructed surfaces on real-world indoor and outdoor scenes from the TNT [22] and Mip-NeRF 360 [2] datasets. Compared to baseline methods, our approach produces more complete and continuous meshes with well-preserved high-frequency details. It also effectively avoids local minima, maintaining fine structures such as holes and sharp edges. By contrast, 2DGS [16] and GOF [56] often yield non-manifold meshes with broken topology, while GS-Pull [60], which extracts surfaces from a learned SDF, tends to produce overly smooth geometry.
Additional comparisons of surface reconstruction on the DTU dataset [18] are shown in Fig. 8. Given the relatively simple geometry of the target objects, most methods achieve visually complete reconstructions. However, the differences lie in the surface quality and detail. Our method demonstrates superior performance in handling reflective surfaces and preserves fine-scale features, producing reconstructions that are both smoother and more accurate, thereby improving visual fidelity and geometric realism.
@@
3DGS [21] 28.94 0.896 0.248 29.93 0.901 0.244 29.43 0.898 0.246
SuGaR [14] 28.71 0.889 0.273 30.12 0.898 0.261 29.41 0.893 0.267
2DGS [16] 28.89 0.898 0.259 29.88 0.901 0.259 29.38 0.899 0.259
GS-Pull [60] 25.69 0.830 0.387 25.89 0.838 0.375 25.79 0.834 0.381
GOF [56] 27.85 0.893 0.257 30.16 0.904 0.242 29.01 0.899 0.250
RaDe-GS [57] 27.83 0.896 0.257 30.04 0.905 0.243 28.94 0.901 0.250
PGSR [4] 28.61 0.891 0.251 29.92 0.903 0.243 29.27 0.897 0.247
@@
3DGS [21] 25.17 21.45 27.18 26.56 22.30 31.34 28.89 30.71 31.98 27.29
2DGS [16] 24.71 21.03 26.61 26.11 22.27 30.72 28.08 30.30 31.24 26.79
GS-Pull [60] 24.19 20.56 26.08 25.24 22.60 30.84 26.41 26.06 29.10 25.68
RaDe-GS [57] 25.56 21.67 27.39 27.10 22.34 30.84 28.73 31.26 31.73 27.40
GOF [56] 25.44 21.59 27.27 26.93 22.40 30.42 28.62 30.64 31.50 27.20
PGSR [4] 25.66 21.52 27.49 26.98 22.29 30.06 28.31 30.80 31.55 27.18
@@
3DGS [21] 0.762 0.602 0.861 0.770 0.631 0.916 0.905 0.923 0.938 0.812
2DGS [16] 0.729 0.569 0.838 0.753 0.614 0.906 0.892 0.915 0.929 0.794
GS-Pull [60] 0.660 0.498 0.776 0.661 0.617 0.896 0.846 0.838 0.897 0.743
RaDe-GS [57] 0.793 0.640 0.870 0.801 0.648 0.916 0.905 0.924 0.939 0.826
GOF [56] 0.786 0.634 0.864 0.790 0.641 0.911 0.900 0.915 0.935 0.820
PGSR [4] 0.793 0.635 0.873 0.798 0.660 0.927 0.914 0.932 0.945 0.831
@@
3DGS [21] 0.216 0.341 0.115 0.219 0.328 0.223 0.204 0.130 0.208 0.220
2DGS [16] 0.276 0.380 0.149 0.263 0.381 0.243 0.230 0.146 0.228 0.255
GS-Pull [60] 0.331 0.443 0.223 0.362 0.414 0.239 0.269 0.248 0.242 0.308
RaDe-GS [57] 0.176 0.286 0.103 0.190 0.279 0.218 0.205 0.130 0.204 0.199
GOF [56] 0.182 0.282 0.109 0.197 0.282 0.221 0.205 0.137 0.200 0.202
PGSR [4] 0.186 0.264 0.103 0.192 0.273 0.180 0.172 0.113 0.169 0.184

### v2

@@ 5 Experiments
Evaluation Protocols. We evaluate our surface reconstruction performance on the DTU [jensen2014large] and Tanks and Temples (TNT) [knapitsch2017tanks] datasets. Following prior works [huang20242d, yu2024gaussian, chen2024pgsr, zhang2024rade], we use 15 scenes from the DTU dataset and 6 scenes from the TNT dataset for evaluation. Depth maps are rendered for all training views, and a TSDF [curless1996volumetric] is constructed for mesh extraction. For novel view synthesis, we use the Mip-NeRF 360 dataset [barron2022mip], which contains large-scale indoor and outdoor scenes with complex lighting and fine-grained geometric details. Following 3DGS [kerbl20233d], one out of every eight images is used for evaluation, while the remaining seven are used for training. We employ COLMAP [schoenberger2016sfm] to generate a sparse point cloud from the original dataset images for initializing the 3D Gaussians. All images are downsampled to a lower resolution to facilitate training. Following established protocols [huang20242d, yu2024gaussian, chen2024pgsr, zhang2024rade], we report Chamfer distance for surface reconstruction on the DTU dataset and F1-score for the TNT dataset. For novel view synthesis, we evaluate using three widely adopted image quality metrics: PSNR, SSIM, and LPIPS.
Implementation Details. Our overall pipeline, training strategy, and hyperparameter settings generally follow 3DGS [kerbl20233d]. We set the number of source views to N=3N\!=\!3, the threshold in $\mathcal{L}_{ns}$ to $\tau\!=\!0.01$, and the patch size in $\mathcal{L}_{p}$ to $7\!\times\!7$. The loss weight factors are set as follows: $\beta_{1}\!=\!0.2$, $\beta_{2}\!=\!0.03$, $\lambda_{1}\!=\!0.015$, $\lambda_{2}\!=\!0.3$, $\lambda_{3}\!=\!0.15$, and $\lambda_{4}\!=\!1.0$. The model is trained for 20,000 iterations for surface reconstruction and 30,000 iterations for novel view synthesis. We first pretrain the model using only the color loss for 7,000 steps to obtain a coarse geometric initialization, which provides a stable foundation for subsequent geometry refinement. Then, we incorporate our image edge item and normal-based geometry alignment into the training. To further refine geometry, we sequentially apply our multi-view photometric alignment for 8,000 iterations, followed by 5,000 iterations of multi-view feature alignment. For novel view synthesis, we continue training for an additional 10,000 steps to optimize rendering quality. All experiments are conducted on a single NVIDIA RTX 4090 GPU.
@@ 5.1 Performance Evaluation
Comparisons on DTU. We first compare our method with state-of-the-art implicit and explicit surface reconstruction approaches on the DTU dataset [jensen2014large]. Following standard protocol, reconstructions are clipped using the provided mask, and evaluations are performed only on foreground objects, as the ground truth point clouds exclude background regions. As shown in Table 1, our method achieves the lowest average Chamfer distance and ranks best across most scenes. Compared to implicit approaches such as NeuS [wang2021neus] and Neuralangelo [li2023neuralangelo], our method delivers significantly better reconstruction accuracy while being much more efficient in terms of runtime. It is worth noting that most implicit methods [wang2021neus, li2023neuralangelo] only reconstruct foreground geometry, whereas our approach can produce detailed and complete meshes, including background regions, which is an essential feature for mesh-based rendering. Although our method is slightly slower than 3DGS [kerbl20233d] and 2DGS [huang20242d] due to the use of multi-view alignment, it achieves significant improvements in reconstruction quality over these earlier Gaussian-based methods.
@@
Comparisons on TNT. We further evaluate our method on the TNT dataset [knapitsch2017tanks], comparing it against both implicit and explicit surface reconstruction baselines. Since the ground-truth point clouds do not include background regions, the evaluation is restricted to foreground objects. As shown in Table 2, our method achieves the best reconstruction performance among all competing approaches, including both implicit and explicit methods. Notably, while several Gaussian-based methods require less optimization time, they tend to produce results with much lower accuracy. In contrast, our method reaches a better balance between efficiency and reconstruction quality. For example, GS-Pull [zhang2024neural] only reconstructs foreground objects and often generates overly smooth surfaces. Fig. 3 provides a qualitative comparison. Our method produces more accurate and detailed reconstructions for both foreground and background regions. It also effectively mitigates the impact of shadows, whereas baseline methods often yield noisy meshes or fail to capture geometric details. The use of geometry, photometric, and feature-based alignment from multiple views provides strong guidance, enabling the Gaussian primitives to converge more accurately to the true surface geometry.
@@
Comparisons on Mip-NeRF 360. We also evaluate our approach on the Mip-NeRF 360 dataset [barron2022mip] for novel view synthesis. Table 3 reports quantitative comparisons against state-of-the-art Gaussian-based and other neural rendering baselines. Our method outperforms competitors on most metrics, demonstrating superior image fitting and generalization to unseen viewpoints. This evidences that our enhanced geometry representation yields higher visual fidelity. Notably, the Mip-NeRF 360 itself achieves the highest average PSNR on indoor scenes but lags on SSIM and LPIPS. Among Gaussian-based methods, 2DGS [huang20242d], SuGaR [guedon2024sugar], and GS-Pull [zhang2024neural] perform worse than vanilla 3DGS [kerbl20233d], suggesting that their planar Gaussian constraints degrade performance in complex environments. Our ablation results in Table 4 further confirm that flattening 3D Gaussians into planar Gaussian disks is ineffective for our framework. Our method preserves the full 3D Gaussian representation and delivers high-quality surfaces without sacrificing novel-view rendering quality. Fig. 4 provides a qualitative comparison of reconstructed meshes. Consistent with our observations on the TNT dataset, our method recovers more accurate and complete surfaces in both foreground and background regions, whereas other methods suffer from noise, oversmoothing, or missing details, especially in challenging indoor scenes.
@@ 5.2 Ablation Studies
(6) Scale regularization: The scaling matrix $\bm{S}$ represents the stretching of a spherical Gaussian along the three axes. Different from previous works [chen2024pgsr, cheng2024gaussianpro, zhang2024neural], incorporating the widely used scale penalty into our method to flatten the 3D Gaussian disks provides no performance gains, and even degrades novel-view rendering quality on the Mip-NeRF 360 dataset.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China (62402401), the Sichuan Provincial Natural Science Foundation of China (2025ZNSFSC1462) and the Fundamental Research Funds for the Central Universities (2682025CX109).
This supplementary document is organized as follows: (1) We first present an overview of the baseline methods used in our evaluation experiments. (2) We then provide additional qualitative and quantitative experimental results to complement those in the main paper. (3) Further ablation studies are included to analyze the impact of key components in our method. (4) More discussion is provided to give a precise understanding of our approach. (5) We discuss the current limitations of our approach and outline potential directions for future research. (6) Finally, we consider the potential negative societal impacts associated with our work.
@@ Appendix A Baseline Methods
We compare our approach with state-of-the-art surface reconstruction and novel view synthesis methods based on 3D Gaussian Splatting, such as 3DGS [kerbl20233d], 2DGS [huang20242d], GOF [yu2024gaussian], and PGSR [chen2024pgsr]. In addition, we evaluate against implicit NeRF-based methods, such as NeuS [wang2021neus] and Neuralangelo [li2023neuralangelo], which utilize Signed Distance Functions (SDFs) to represent scenes and convert them into opacity fields for volume rendering via ray tracing.
@@ B.1 Comparisons on the Deep Blending Dataset
In order to further test the performance of the algorithm on different real-world data, we evaluate both novel view synthesis and surface reconstruction on two commonly used indoor scenes, Dr Johnson and Playroom, from the Deep Blending dataset [hedman2018deep]. Table 5 presents quantitative comparisons against state-of-the-art Gaussian-based methods on novel view synthesis. All results, unless otherwise noted, are obtained from our own re-implementations using the official code released by the respective authors. For 3DGS [kerbl20233d], we use the pretrained models provided by the authors. For SuGaR [guedon2024sugar], we report numbers directly from the original paper to avoid potential inconsistencies. Our method consistently outperforms all baselines across all metrics, demonstrating superior rendering quality and generalization to unseen viewpoints. These improvements can be attributed to our enhanced geometric representation, which yields better visual fidelity. Consistent with observations in the evaluation on the Mip-NeRF 360 dataset [barron2022mip] presented in the main paper, methods such as 2DGS [huang20242d], SuGaR [guedon2024sugar], and GS-Pull [zhang2024neural] underperform compared to vanilla 3DGS [kerbl20233d], suggesting that imposing planar Gaussian constraints may hinder performance in complex scenes.
@@ B.2 More Experimental Results
Novel view synthesis. To provide a more comprehensive evaluation of novel view synthesis, we report per-scene quality metrics (PSNR, SSIM, and LPIPS) for Gaussian-based methods on the Mip-NeRF 360 dataset [barron2022mip], as shown in Table 6. All results are obtained from our own runs using the official implementations of prior methods, except for 3DGS, where we use the pretrained models provided by the authors to avoid training inconsistencies. Our method consistently achieves the highest PSNR and SSIM scores on most scenes and has the best average scores. Notably, it yields significant improvements in LPIPS across all scenes, highlighting its ability to capture high-frequency details and produce perceptually superior renderings.
@@
Surface reconstruction. Fig. 7 presents additional qualitative comparisons of reconstructed surfaces on real-world indoor and outdoor scenes from the TNT [knapitsch2017tanks] and Mip-NeRF 360 [barron2022mip] datasets. Compared to baseline methods, our approach produces more complete and continuous meshes with well-preserved high-frequency details. It also effectively avoids local minima, maintaining fine structures such as holes and sharp edges. By contrast, 2DGS [huang20242d] and GOF [yu2024gaussian] often yield non-manifold meshes with broken topology, while GS-Pull [zhang2024neural], which extracts surfaces from a learned SDF, tends to produce overly smooth geometry.
Additional comparisons of surface reconstruction on the DTU dataset [jensen2014large] are shown in Fig. 8. Given the relatively simple geometry of the target objects, most methods achieve visually complete reconstructions. However, the differences lie in the surface quality and detail. Our method demonstrates superior performance in handling reflective surfaces and preserves fine-scale features, producing reconstructions that are both smoother and more accurate, thereby improving visual fidelity and geometric realism.
@@
3DGS [kerbl20233d] 28.94 0.896 0.248 29.93 0.901 0.244 29.43 0.898 0.246
SuGaR [guedon2024sugar] 28.71 0.889 0.273 30.12 0.898 0.261 29.41 0.893 0.267
2DGS [huang20242d] 28.89 0.898 0.259 29.88 0.901 0.259 29.38 0.899 0.259
GS-Pull [zhang2024neural] 25.69 0.830 0.387 25.89 0.838 0.375 25.79 0.834 0.381
GOF [yu2024gaussian] 27.85 0.893 0.257 30.16 0.904 0.242 29.01 0.899 0.250
RaDe-GS [zhang2024rade] 27.83 0.896 0.257 30.04 0.905 0.243 28.94 0.901 0.250
PGSR [chen2024pgsr] 28.61 0.891 0.251 29.92 0.903 0.243 29.27 0.897 0.247
@@
3DGS [kerbl20233d] 25.17 21.45 27.18 26.56 22.30 31.34 28.89 30.71 31.98 27.29
2DGS [huang20242d] 24.71 21.03 26.61 26.11 22.27 30.72 28.08 30.30 31.24 26.79
GS-Pull [zhang2024neural] 24.19 20.56 26.08 25.24 22.60 30.84 26.41 26.06 29.10 25.68
RaDe-GS [zhang2024rade] 25.56 21.67 27.39 27.10 22.34 30.84 28.73 31.26 31.73 27.40
GOF [yu2024gaussian] 25.44 21.59 27.27 26.93 22.40 30.42 28.62 30.64 31.50 27.20
PGSR [chen2024pgsr] 25.66 21.52 27.49 26.98 22.29 30.06 28.31 30.80 31.55 27.18
@@
3DGS [kerbl20233d] 0.762 0.602 0.861 0.770 0.631 0.916 0.905 0.923 0.938 0.812
2DGS [huang20242d] 0.729 0.569 0.838 0.753 0.614 0.906 0.892 0.915 0.929 0.794
GS-Pull [zhang2024neural] 0.660 0.498 0.776 0.661 0.617 0.896 0.846 0.838 0.897 0.743
RaDe-GS [zhang2024rade] 0.793 0.640 0.870 0.801 0.648 0.916 0.905 0.924 0.939 0.826
GOF [yu2024gaussian] 0.786 0.634 0.864 0.790 0.641 0.911 0.900 0.915 0.935 0.820
PGSR [chen2024pgsr] 0.793 0.635 0.873 0.798 0.660 0.927 0.914 0.932 0.945 0.831
@@
3DGS [kerbl20233d] 0.216 0.341 0.115 0.219 0.328 0.223 0.204 0.130 0.208 0.220
2DGS [huang20242d] 0.276 0.380 0.149 0.263 0.381 0.243 0.230 0.146 0.228 0.255
GS-Pull [zhang2024neural] 0.331 0.443 0.223 0.362 0.414 0.239 0.269 0.248 0.242 0.308
RaDe-GS [zhang2024rade] 0.176 0.286 0.103 0.190 0.279 0.218 0.205 0.130 0.204 0.199
GOF [yu2024gaussian] 0.182 0.282 0.109 0.197 0.282 0.221 0.205 0.137 0.200 0.202
PGSR [chen2024pgsr] 0.186 0.264 0.103 0.192 0.273 0.180 0.172 0.113 0.169 0.184